OpenStudy (kaylala):

Trigonometric Equation 2cos^2A - cosA = 0

OpenStudy (amoodarya):

factor cos A cos A (2cosA -1)=0 or use substitution cos A = x so 2x^2-x=0

OpenStudy (kaylala):

Then what? @amoodarya

OpenStudy (kaylala):

@Callisto

OpenStudy (callisto):

Do you know how to solve a quadratic equation?

OpenStudy (kaylala):

Yes. But i dont know how to do this?

OpenStudy (callisto):

Let x = cosA 2cos^2A - cosA = 0 is then rewritten as 2x^2 - x = 0 First, solve for x. What do you get?

OpenStudy (kaylala):

I got: x(2x-1) Which will be x=0, x=1/2

OpenStudy (callisto):

Substitute x=cosA again into the solutions you got, i.e. cosA = 0 or cosA = 1/2 Can you solve A now?

OpenStudy (kaylala):

What formula shall i use? I'm really sorry i' new to this

OpenStudy (kaylala):

Wait let me try

OpenStudy (callisto):

You need not use formula. Just a table and a graph.

OpenStudy (kaylala):

Isee Cos A=0 A={0,180} Cos A=1/2 A={60,300} Did i do it correctly???

OpenStudy (callisto):